OpenStudy (anonymous):

Factor completely. 4x^2+8xy+4y^2

OpenStudy (anonymous):

perfect square

OpenStudy (anonymous):

4(x^2+2xy+y^2)=4(x^2+xy+xy+y^2)=4(x+y)^2